40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 7334     Entry time: Tue Sep 4 11:32:58 2012
Author: Jenne 
Type: Update 
Category: Locking 
Subject: Friday in-vac work 

Elog re: Friday's work

Adjusted PZT2 so we're hitting the center of PR2. 

Noticed that the beam centering target is too low by a few mm, since the OSEM set screw holes that it mounts to are lower than the center line of the optic.  This meant that while we were hitting the center of PR2, the beam was half clipped by PRM's centering target.  We removed the target to confirm that the beam is really centered on PR2.

Checked the beam on PR3 - it looked fine.  There had been concern last week that PR2 was severely pitched forward, but this turns out to be an effect of the PRM centering target being too low - shoot the beam downward to go through the hole, beam continues downward to hit the bottom of PR2, so beam is falling of the bottom of PR3.  But when we actually centered the beam on PR2, things looked fine on PR3. 

Checked that the beam approximately goes through the beam splitter.  Again, the targets are too low, and these 45 deg targets' holes are smaller than the 0 deg targets, so we don't see any beam going through the target, since the beam is hitting the target higher than the hole.  The beam looked left/right like it was pretty close to the hole, but it was hard to tell since the angle is bad, and I'm not infinitely tall.  We should check again to make sure that the beam is going through properly, and we're not clipping anywhere.  I'll need help from a height-advantaged person for this.

Checked that the beam is hitting the center of the ITMY, as best we can see by using an IR card at the back of the optic.  We didn't try reaching around to put a target on the front side. 

We were debating whether it would be worth it to open ETMY this week, to check that the beam transmitted through the BS hits the center of ETMY.

We also took a quick look around the AS optics, but since that depends on BS/ITMX alignment, we weren't sure how to proceed.  We need a plan for this part.  All suspended optics were restored to their last good alignment, but we haven't tried locking MICH or anything to confirm that the alignment. 

To do list:  Check no clipping on ITMY table of beam between BS and ITMY, clipping on POY optics.  Also, oplev is clipping on cable holder thing on the table - this needs to be moved.  .....other?

ELOG V3.1.3-